# WER We Stand: Benchmarking Urdu ASR Models

### v1

@@ Abstract
Samee Arif1, Aamina Jamal Khan1*, Mustafa Abbas1*, Agha Ali Raza1, Awais Athar2, {samee.arif, 25100162, 25100079, agha.ali.raza}@lums.edu.pk, awais@ebi.ac.uk 1Lahore University of Management Sciences, 2EMBL European Bioinformatics Institute
@@ 1 Introduction
The code, and model outputs are publicly available on GitHub [^1] https://github.com/ulrs0/Urdu-ASR-Today.
@@ 2.2 Low-Resource ASR
The ASRoIL survey (Singh et al., 2020) provides valuable insights for Indian languages into the challenges of variability in speech signals and the availability of corpora, detailing feature extraction techniques like MFCC, LPCC, and PLP, as well as various modeling and classification techniques. Unnibhavi and Jangamshetti, 2016’s work offers an overview of ASR systems for South Indian languages, explaining methodologies and feature extraction methods by giving digests of seven papers. Dhouib et al., 2022’s work systematically reviewed Arabic ASR studies from 2011 to 2021, covering feature extraction and classification techniques along with various aspects, such as the types of Arabic language supported, performance metrics, and existing research gaps. Abdelhamid et al., 2020 focus on end-to-end deep learning frameworks, which integrate and simultaneously train the language model, pronunciation, and acoustic components, thus simplifying the pipeline. Kadyan et al., 2019 presents a comparative study of Deep Neural Network (DNN) based Punjabi-ASR systems demonstrating that DNN-HMM hybrid models outperform traditional GMM-HMM architectures.
Low-resource languages have demonstrated significant advancements in multilingual settings, where models benefit from unexpected but advantageous learning capabilities. Toshniwal et al., 2018 found that multilingual models exhibit cross-lingual transfer learning, which enhances the accuracy of low-resource languages by leveraging similarities and commonalities from high-resource languages. Tüske et al., 2013 further analyzed cross-lingual transfer effects, finding that multilingual models fine-tuned on low-resource languages outperformed monolingual models, demonstrating the effectiveness of cross-lingual transfer. Overall, their work shows that low-resource languages like Urdu are more sensitive to model architecture due to data scarcity, and multilingual models can be trained effectively for both low-resource multilingual and monolingual tasks.
@@ 2.3 Urdu ASR
To the best of our knowledge, there is no up-to-date survey on the state of ASR in Urdu. Ashraf et al., 2010 developed a small-sized GMM using Word List Grammar by processing individual words and combining it with a Knowledge Base storing audio representations, pronunciations, and probabilities, achieving a Word Error Rate (WER) of 5.33 on seen speakers and 10.66 on unseen speakers. Asadullah et al., 2016 explored the impact of vocabulary size on HMM performance for Urdu ASR, finding an insignificant WER increase of 0.5% between 100 and 250-word vocabularies. Naeem et al., 2020 used a GMM combined with a graphene-to-phone LSTM CMUsphinx and KENLM to calculate and store n-gram probabilities, achieving a 9.64% WER, significantly improving accuracy. Khan et al., 2023 combined HMM and CNN-TDNN with LF-MMI, achieving a 13% WER.
@@ 5.1.1 Quantitative Analysis
Several words stand out as high-frequency errors across all models. Notably, "Ho ga" and "Chahiye" are consistently problematic, with particularly high error counts in whisper-medium, whisper-large, and seamless-large. The word "Chahiye" misspelled in Urdu: one with two "Yeh" and a "Hamza" characters and one with one "Yeh" and a "Hamza". These errors can likely be attributed to the training data, which often includes user-generated content such as internet text, where such misspellings are common. This noisy data likely contributes to the model’s tendency to replicate these errors.
"Chahiye" (correct one) has an error frequency of 1,269 in whisper-medium and 3,173 in whisper-large, indicating that these models struggle significantly with this specific word. mms-1b shows relatively fewer errors for many words in this specific analysis, but it has a higher WER for the ARL and CSaLT datasets. This indicates that, while fewer individual words may be highlighted as problematic here, the model struggles with a broader range of words overall, leading to a higher overall error rate . The word "Hoga" shows a high error rate across all models, as it should be written as "Ho ga" in the model transcriptions, but the test dataset incorrectly contains it as a single word, "Hoga." This mismatch leads to consistent false positives, where the models’ correct transcription is flagged as incorrect due to the way the word is represented in the test dataset. The Whisper family has the higher tendency to misrecognize the word "Hai" followed by MMS and then Seamless.
Figure 3 compares the most frequent errors on ARL and CSaLT datasets across the fine-tuned models. In the fine-tuned models, certain high-frequency errors observed in the non-fine-tuned models have been reduced, indicating that fine-tuning has helped address some transcription challenges. However in some cases the error rate has gone up. For instance, the error rate for the word "Chahiye" (correct one) has increased after fine-tuning. We found 155 misspelled versions of this word in Common Voice and Fleurs dataset. For seamless-medium, the error has gone up from 0 to 1942. whisper-large still shows a high error with slight increase from 3173 to 3391. The word "Ho ga" also poses a challenge for fine-tuned models. There are 82 instances of "Ho ga" (correct) and 87 instances of "Hoga" (incorrect) in the training dataset. As a result, the false positives seen earlier are now turned into inconsistencies, where the models are confused between the two variations, leading to unpredictable outputs.
Figure 3: Comparison of wrong words across fine-tuned models.
@@ 5.2.1 Quantitative Analysis
Figure 4: Comparison of substitution errors across models before and after fine-tuning.
Figure 5: Comparison of deletion errors across models before and after fine-tuning.
@@ 5.2.2 Qualitative Analysis
A significant limitation across all ASR models is their inability to handle overlapping speech from multiple speakers. As illustrated in Figure 6, the models tend to transcribe only the dominant speaker, ignoring other concurrent voices.
Figure 6: Example of multi-speaker audio.
@@
Figure 7: Translation and paraphrasing error in Seamless M4T. The yellow box represents the ground truth and the blue ones represents the prediction.
@@ 7 Future Work
This study opens several avenues for future exploration in Urdu ASR. Addressing overlapping speech in conversational settings remains a critical challenge. Future work could focus on evaluating speaker diarization using our dataset to improve multi-speaker recognition. Enhancing the handling of code-switching between Urdu and English, especially for models like seamless-large that tend to translate rather than transliterate, could significantly boost transcription accuracy in real-world contexts where multilingualism is common. Exploring the potential of Large Language Models (LLMs), such as GPT-4o and Llama-3.1, for ASR output post-processing could further refine transcription quality by correcting errors and improving fluency. Finally, the integration of a robust text normalization system tailored for Urdu could address variations in spelling and word forms, improving the consistency of transcriptions. Extending this work to cross-modal tasks, such as speech-to-speech translation or text generation, could provide valuable insights for building more advanced systems that handle complex, multimodal language tasks for low-resource languages like Urdu.
@@ Acknowledgments
We are grateful to our research interns: Muhammad Abubakar Mughal, Natiq Khan, Anum Javed, Aleena Saqib, Muhammad Abdullah Sohail, Bushra Zubair, Hamza Iqbal, Salaar Masood, Maham Javed, Faisal Haider, Muhammad Suhaib Rashid, and Muhammad Faizan Waris. Their dedication and hard work in creating and refining the dataset were instrumental to the success of this project. We sincerely appreciate their invaluable contributions and commitment to our research efforts.

### v3

@@ Abstract
Samee Arif1, Sualeha Farid2, Aamina Jamal Khan1*, Mustafa Abbas1*, Agha Ali Raza1, Awais Athar3, {samee.arif, 25100162, 25100079, agha.ali.raza}@lums.edu.pk, sualeha@umich.edu, awais@ebi.ac.uk, 1Lahore University of Management Sciences, 2University of Michigan - Ann Arbor, 3EMBL European Bioinformatics Institute
@@ 1 Introduction
The code, and model outputs are publicly available on GitHub [^1] https://github.com/sameearif/WER-We-Stand.
@@ 2.2 Low-Resource ASR
The ASRoIL survey (Singh et al., 2020) provides valuable insights for Indian languages into the challenges of variability in speech signals and the availability of corpora, detailing feature extraction techniques like MFCC, LPCC, and PLP, as well as various modeling and classification techniques. Unnibhavi and Jangamshetti (2016)’s work offers an overview of ASR systems for South Indian languages, explaining methodologies and feature extraction methods by giving digests of seven papers. Javed et al. (2022) introduce the IndicSUPERB benchmark, featuring 1,684 hours of labeled speech data and establish benchmarks for six speech tasks. They train and evaluate different self-supervised models and demonstrate that language-specific fine-tuned models outperform baseline methods. Dhouib et al. (2022)’s work systematically reviewed Arabic ASR studies from 2011 to 2021, covering feature extraction and classification techniques along with various aspects, such as the types of Arabic language supported, performance metrics, and existing research gaps. Abdelhamid et al. (2020) focus on end-to-end deep learning frameworks, which integrate and simultaneously train the language model, pronunciation, and acoustic components, thus simplifying the pipeline. Kadyan et al. (2019) presents a comparative study of Deep Neural Network (DNN) based Punjabi-ASR systems demonstrating that DNN-HMM hybrid models outperform traditional GMM-HMM architectures.
Low-resource languages have demonstrated significant advancements in multilingual settings, where models benefit from unexpected but advantageous learning capabilities. Toshniwal et al. (2018) found that multilingual models exhibit cross-lingual transfer learning, which enhances the accuracy of low-resource languages by leveraging similarities and commonalities from high-resource languages. Tüske et al. (2013) further analyzed cross-lingual transfer effects, finding that multilingual models fine-tuned on low-resource languages outperformed monolingual models, demonstrating the effectiveness of cross-lingual transfer. Overall, their work shows that low-resource languages like Urdu are more sensitive to model architecture due to data scarcity, and multilingual models can be trained effectively for both low-resource multilingual and monolingual tasks.
@@ 2.3 Urdu ASR
Sharif et al. (2024) provide an amalgamation of overviews of relevant studies on Urdu ASR, highlighting current trends, technological advancements, and future research directions, without directly comparing results on a common dataset. Farooq et al. (2019) develop an Urdu LVCSR system using 300 hours of speech data and explores various acoustic modeling techniques, achieving a WER of 13.50%, though the models and data are not publicly available. Ashraf et al. (2010) developed a small-sized GMM using Word List Grammar by processing individual words and combining it with a Knowledge Base storing audio representations, pronunciations, and probabilities, achieving a Word Error Rate (WER) of 5.33 on seen speakers and 10.66 on unseen speakers. Asadullah et al. (2016) explored the impact of vocabulary size on HMM performance for Urdu ASR, finding an insignificant WER increase of 0.5% between 100 and 250-word vocabularies. Naeem et al. (2020) used a GMM combined with a graphene-to-phone LSTM CMUsphinx and KENLM to calculate and store n-gram probabilities, achieving a 9.64% WER, significantly improving accuracy. Khan et al. (2023) combined HMM and CNN-TDNN with LF-MMI, achieving a 13% WER.
@@ 5.1.1 Quantitative Analysis
Several words appear to stand out as high-frequency errors across all models. Notably, "Ho ga" and "Chahiye" are consistently problematic, with particularly high error counts in whisper-medium, whisper-large, and seamless-large. The word "Chahiye" appears with orthographic variations in Urdu—one with two "Yeh" and no "Hamza" character, one with one "Yeh" and a "Hamza," and one with with no "Yeh" and a "Hamza." The first variation is the most widely accepted correct form, though all are used interchangeably in practice. These errors can likely be attributed to the training data, which often includes user-generated content like internet text, where such variations are common.
"Chahiye" has an error frequency of 1,269 in whisper-medium and 3,173 in whisper-large, indicating that these models do not reproduce the specific spelling of this word as used in the dataset. mms-1b shows relatively fewer errors for many words in this specific analysis, but it has a higher WER for the ARL and CSaLT datasets. This indicates that, while fewer individual words may be highlighted as problematic here, the model struggles with a broader range of words overall, leading to a higher overall error rate. The word "Hoga" shows a high error rate across all models, as it is reproduced as "Ho ga" in the model transcriptions, but the test dataset contains it as a single word, "Hoga." This mismatch leads to consistent false positives, where the models’ correct transcription is flagged as incorrect due to the way the word is represented in the test dataset. The Whisper family has the higher tendency to misrecognize the word "Hai" followed by MMS and then Seamless.
Figure 3 compares the most frequent errors on ARL and CSaLT datasets across the fine-tuned models. In the fine-tuned models, certain high-frequency errors observed in the non-fine-tuned models have been reduced, indicating that fine-tuning has helped address some transcription challenges. However in some cases the error rate has gone up. For instance, the error rate for the word "Chahiye" (expected spellings) has increased after fine-tuning. We found 155 misspelled versions of this word in Common Voice and Fleurs dataset. For seamless-medium, the error has gone up from 0 to 1942. whisper-large still shows a high error with slight increase from 3173 to 3391. The word "Ho ga" also poses a challenge for fine-tuned models. There are 82 instances of "Ho ga" and 87 instances of "Hoga" in the training dataset. As a result, the false positives seen earlier are now turned into inconsistencies, where the models are confused between the two variations, leading to unpredictable outputs. A lattice of plausible respellings of the reference transcription, as done so by Karita et al., 2023 may help address this issue. By employing a lenient evaluation with orthographical variations in mind, these false errors could be overlooked.
Figure 3: Wrong words across fine-tuned models.
@@ 5.2.1 Quantitative Analysis
Figure 4: Substitution errors across models before and after fine-tuning.
Figure 5: Deletion errors across models before and after fine-tuning.
@@ 5.2.2 Qualitative Analysis
A significant limitation across all ASR models is their inability to handle overlapping speech from multiple speakers. Figure 6 illustrates an example of ground truth for overlapping speech, where the audio from speaker 1 and speaker 2 overlaps. As shown in the figure, speaker 1’s audio is dominant, causing ASR models to fail in accurately transcribing the speech of speaker 2.
Figure 6: The blue color represents speaker 1 and yellow represents speaker 2.
@@
Figure 7: The yellow box represents the ground truth and the blue ones represents the prediction.
@@ 7 Future Work
This study opens several avenues for future exploration in Urdu ASR. Addressing overlapping speech in conversational settings remains a critical challenge. Future work could focus on evaluating speaker diarization using our dataset to improve multi-speaker recognition. Enhancing the handling of code-switching between Urdu and English, especially for models like seamless-large that tend to translate rather than transliterate, could significantly boost transcription accuracy in real-world contexts where multilingualism is common. Exploring the potential of Large Language Models (LLMs), such as GPT-4o and Llama-3.1, for ASR output post-processing could further refine transcription quality by correcting errors and improving fluency. Finally, the integration of a lenient evaluation method which incorporates orthographical variations, or robust text normalization system tailored for Urdu could address variations in spelling and word forms, improving the consistency of transcriptions. Extending this work to cross-modal tasks, such as speech-to-speech translation or text generation, could provide valuable insights for building more advanced systems that handle complex, multimodal language tasks for low-resource languages like Urdu.
@@ Acknowledgments
We are deeply grateful to Sualeha Farid (University of Michigan - Ann Arbor) for her invaluable contributions, including conducting data generation sessions and assisting with data analysis, without which this paper would not have been possible.
We are also grateful to our research interns: Muhammad Abubakar Mughal, Natiq Khan, Anum Javed, Aleena Saqib, Muhammad Abdullah Sohail, Bushra Zubair, Hamza Iqbal, Salaar Masood, Maham Javed, Faisal Haider, Muhammad Suhaib Rashid, and Muhammad Faizan Waris. Their dedication and hard work in creating and refining the dataset were instrumental to the success of this project. We sincerely appreciate their invaluable contributions and commitment to our research efforts.